OpenStudy (anonymous):

John made a rectangle pen for his dog using 28 feet of fencing. If the width of the pen is 2 feet more than one-half the length, what is the length and width of the pen? The length of the pen is 6 feet 7 feet 8 feet

OpenStudy (ajprincess):

Let l be the length and w be the width. w=l/2+2 Perimeter of the rectangle pen=Total length of fencing =28 =2(l+w) 2(l+w)=28 2(l+(l/2)+2)=28 Can u find l nw? @806

OpenStudy (anonymous):

hi 806 get on chat